# AutoFT: Robust Fine-Tuning by Optimizing Hyperparameters on OOD Data

### 18 Jan 2024 v1

@@ Abstract
Foundation models encode rich representations that can be adapted to a desired task by fine-tuning on task-specific data. However, fine-tuning a model on one particular data distribution often compromises the model’s original performance on other distributions. Current methods for robust fine-tuning utilize hand-crafted regularization techniques to constrain the fine-tuning process towards the base foundation model. Yet, it is hard to precisely specify what characteristics of the foundation model to retain during fine-tuning, as this depends on how the pre-training, fine-tuning, and evaluation data distributions relate to each other. We propose AutoFT, a data-driven approach for guiding foundation model fine-tuning. AutoFT optimizes fine-tuning hyperparameters to maximize performance on a small out-of-distribution (OOD) validation set. To guide fine-tuning in a granular way, AutoFT searches a highly expressive hyperparameter space that includes weight coefficients for many different losses, in addition to learning rate and weight decay values. We evaluate AutoFT on nine natural distribution shifts which include domain shifts and subpopulation shifts. Our experiments show that AutoFT significantly improves generalization to new OOD data, outperforming existing robust fine-tuning methods. Notably, AutoFT achieves new state-of-the-art performance on the WILDS-iWildCam and WILDS-FMoW benchmarks, outperforming the previous best methods by $6.0\%$ and $1.5\%$, respectively.
@@ 1 Introduction
Foundation models have emerged as a powerful tool in machine learning, demonstrating unprecedented performance across a wide variety of data distributions  [72, 40, 41]. By pre-training on large and diverse datasets, these models learn representations that can serve as rich common-sense priors that complement task-specific data. We thus expect fine-tuning to enhance the generalization capabilities of foundation models. However, fine-tuning often degrades the performance of foundation models on out-of-distribution (OOD) data. This indicates that conventional fine-tuning strategies can fail to utilize the prior knowledge embedded in the foundation model.
@@
Figure 1: Overview of AutoFT: a method for robustly fine-tuning foundation models. While fine-tuning with in-distribution (ID) data (blue), AutoFT optimizes hyperparameters for post-adaptation performance on a small out-of-distribution validation set (red). This validation set serves as a proxy for performance on different distributions (green and purple), encouraging AutoFT to learn a robust fine-tuning procedure.
@@
Figure 2: A summary of our data assumptions and evaluation protocol. The standard approach is to optimize hyperparameters on a validation dataset drawn from the same distribution as the training data. In contrast, AutoFT employs a small out-of-distribution (OOD) validataion set for hyperparameter optimization, enhancing the generalizability of the final model. We evaluate all fine-tuned models on data from unseen distribution shifts (green and purple).
@@
This issue of conventional fine-tuning distorting beneficial foundation model priors has driven recent research on developing robust fine-tuning methods. Such methods aim to produce an adapted model that achieves good performance under distribution shifts by preserving the prior knowledge embedded in the foundation model. Prior works have proposed various regularization techniques for this purpose, such as ensembling models before and after adaptation [92] or initially fitting only the last layer [50]. However, as these methods are primarily based on human intuition, they may not fully account for the complex interplay between the foundation model priors and the adaptation process.
We introduce AutoFT, a novel method for robust fine-tuning that aims to find the right tradeoff between the prior and the fine-tuning data through hyperparameter optimization. Our key insight is that we can learn what characteristics of the foundation model to preserve during fine-tuning using a data-driven approach. Like existing robust fine-tuning methods, we fine-tune a foundation model on task-specific data, and then evaluate the resulting model on a set of OOD distributions. However, we additionally leverage a small OOD validation set with up to $1000$ labeled examples from one unseen distribution; we optimize fine-tuning hyperparameters for post-adaptation performance on this OOD validation set. Importantly, the OOD validation set is only used for hyperparameter optimization, not fine-tuning, and does not follow the same distribution as the OOD test sets. We illustrate the intuition behind our approach in Figure 1 and our data assumptions in Figure 2.
We make two key alterations to standard hyperparameter optimization, which we find to be critical for automatic robust fine-tuning. First, as mentioned above, we optimize hyperparameters with respect to an OOD validation set rather than an ID validation set. Second, we use a broader definition of “hyperparameter”: beyond the usual hyperparameters such as learning rate, we parameterize the fine-tuning objective itself through weight coefficients for several different loss functions and regularizers. This larger hyperparameter search space gives AutoFT more granular control over adaptation.
@@
We rigorously evaluate AutoFT on a wide array of real-world datasets and consider various types of distribution shifts, including subpopulation and domain shift. Our experiments show that our approach results in better generalization to unseen OOD data. With at most $1000$ datapoints from an OOD distribution, AutoFT outperforms existing robust fine-tuning methods across all benchmarks. These gains in robustness are achieved with minimal additional compute, requiring at most $5\%$ more total compute compared to standard fine-tuning. Among other results, AutoFT achieves new state-of-the-art performance on the challenging iWildCam and FMoW benchmarks  [9, 47, 16], outperforming the prior best methods by $6.0\%$ and $1.5\%$, respectively.
@@ 2 Related Work
AutoML and hyperparameter optimization. Our work leverages high-level ideas from the broader literature on meta-learning and hyperparameter optimization. Such methods have proposed to optimize different parts of the training pipeline, including general hyperparameters [39, 11, 26, 58, 59], network architectures [100, 101, 62, 74, 93], augmentation policies [19, 60, 34, 20], optimizers [10, 4, 89, 65, 15], and objective functions [97, 46, 68, 8]. However, most of these works optimize for generalization within the training distribution and do not consider robustness to distribution shifts. Existing works that optimize a training procedure for OOD generalization consider a structured few-shot adaptation setting [56, 99], limiting their scalability to large datasets.
@@
Transfer learning. While early research demonstrated the effectiveness of using features learned from pre-training on large datasets for new tasks, transfer learning has evolved to focus on optimizing performance in limited data scenarios [69, 96, 81]. Common transfer learning techniques include fine-tuning with regularization [98, 94, 52, 42, 57, 1, 29, 82, 43] and selectively freezing pre-trained parameters [45, 53, 33, 73, 63, 77, 22, 24, 23, 17, 85, 54, 51]. However, as pre-trained models are fine-tuned for a specific distribution, their effective robustness decreases at convergence [3]. We introduce a transfer learning framework that preserves the robustness of the pre-trained model while adapting to new tasks.
@@
Out-of-distribution generalization. Maintaining good performance on data that deviates from the training distribution is crucial in many real-world applications, where models may face data from unfamiliar environments [35, 28, 32, 47]. Numerous studies have investigated how to ensure robustness to various distribution shifts [86, 14, 36, 5, 80, 61, 90, 3, 18, 55]. Some works have shown that despite the myriad of ways in which data distributions can change, naturally occurring distribution shifts have a surprisingly predictable effect on model performance [83, 66, 6], suggesting that it may be possible to learn how to be robust to unseen natural distribution shifts. Goyal et al. [31] is related to our work: they dynamically adapt models to distribution shifts via a meta-learned test-time adaptation loss applied to unlabeled test data. AutoFT instead focuses on fine-tuning foundation models and does not require any test data during training. AutoFT improves generalization to novel distribution shifts after fine-tuning using only a small OOD validation set.
@@
Robust fine-tuning. Foundation models trained on massive datasets encode a broad range of general knowledge, enabling robust performance across various data distributions, including OOD scenarios [72, 13]. While in principle, foundation models should serve as a useful prior for further fine-tuning, empirical evidence shows that fine-tuning foundation models on a new task often leads to a significant drop in OOD performance [3]. Recent works have proposed modifications to the basic fine-tuning procedure to improve OOD generalization [92, 50, 91, 30, 67]. Instead of hand-designing a regularization technique, we propose a data-driven approach to learn a more nuanced fine-tuning procedure. In fact, some prior works [94, 54, 30] can be seen as special cases of AutoFT since our hyperparameter search space encompasses these fine-tuning algorithms. Our experiments in Sec. 6 demonstrate that AutoFT consistently outperforms prior robust fine-tuning methods on OOD data.
3 Hyperparameter Optimization
We begin by formalizing hyperparameter optimization, a procedure that we extend for robustly fine-tuning foundation models in Sec. 4. Hyperparameters are predefined properties of the learning algorithm which are not learned during training, such as network architecture, learning rate, and regularization strength. Since hyperparameters significantly impact the performance of the final model, it is crucial to choose the right hyperparameters for any learning problem. It is difficult to know the optimal hyperparameters a priori, as they are influenced by many properties of the problem setting, including the data distribution and desired performance metric.
@@ 3 Hyperparameter Optimization
Formally, we denote the learning algorithm as LearnAlg and its hyperparameters as $\phi\in\Phi$, where $\Phi$ is the hyperparameter space. We also denote the training and validation datasets as $D_{\text{tr}}$ and $D_{\text{val}}$, respectively. These datasets are disjoint and are typically drawn from the same distribution. We denote the resulting model as $\text{LearnAlg}(\phi,D_{\text{tr}})$, to explicitly represent the dependence on both hyperparameters $\phi$ and training data $D_{\text{tr}}$. The goal of hyperparameter optimization is to find hyperparameters that maximize some performance metric $\mathrm{Perf}(f,D_{\text{val}})$ which depends on the model $f$ and the validation dataset $D_{\text{val}}$. Examples of performance metrics include top-1 accuracy, macro F1 score, and worst-region accuracy. We can formally state the hyperparameter optimization problem as
@@
Prior hyperparameter optimization methods typically start with randomly initialized model parameters and use a validation set $D_{\text{val}}$, drawn from the same distribution as the training data, to adjust hyperparameters. The problem of robust fine-tuning, however, begins with pre-trained model parameters and aims to achieve high performance on test data that diverges from the original training data’s distribution. In the next section, we describe how we modify the standard hyperparameter optimization loop for robustly fine-tuning pre-trained foundation models.
4 AutoFT: Robust Fine-Tuning via Hyperparameter Optimization
In this section, we present AutoFT, a data-driven method for robustly fine-tuning large pre-trained models. AutoFT aims to find hyperparameters for fine-tuning a pre-trained model that strike a better balance between the general-purpose knowledge encoded in the pre-trained model and the task-specific knowledge in the training data. To achieve this, AutoFT optimizes hyperparameters with respect to a performance metric that is sensitive to degradation under distribution shift. Specifically, we leverage a small validation set from a different distribution from the fine-tuning data; performance on this OOD validation set serves as a proxy for general OOD performance.
@@ 4.1 Problem Setting: Robust Fine-Tuning
During hyperparameter optimization, we assume access to two datasets: (1) a large fine-tuning dataset $D_{\text{tr}}$ from the training distribution $\mathcal{P}_{\text{tr}}$, and (2) a small held-out OOD validation set $D_{\text{val}}$ from a different distribution $\mathcal{P}_{\text{val}}$. The validation set $D_{\text{val}}$ is much smaller than $D_{\text{tr}}$, and is used exclusively for hyperparameter optimization, not for fine-tuning. After hyperparameter optimization, we fine-tune on $D_{\text{tr}}$ with the best found hyperparameters and then test the fine-tuned model on several OOD distributions $\mathcal{P}_{\text{ood}}^{1},\mathcal{P}_{\text{ood}}^{2},\dots$. These test distributions are different from both $\mathcal{P}_{\text{tr}}$ and $\mathcal{P}_{\text{val}}$, and are not shown to the model during either fine-tuning or hyperparameter optimization. We visualize this data assumption in Figure 2.
Let $f$ denote a pre-trained model with parameters $\theta$, which we adapt to the task at hand by fine-tuning on $D_{\text{tr}}$. We consider foundation models $f$ that possess rich, general-purpose knowledge from being pretrained on large, diverse datasets. Let $\phi\in\Phi$ represent the hyperparameters for fine-tuning, and let $\text{LearnAlg}(\phi,D_{\text{tr}})$ denote the fine-tuning algorithm which produces adapted model parameters. In this work, the fine-tuning algorithm LearnAlg is stochastic gradient descent. The performance metric $\mathrm{Perf}$ differs across the experiments we consider: it is top-1 accuracy for most settings.
We assume that all data distributions pertain to the same task, with differences arising from domain shifts caused by natural variation in the data collection process. For example, all data may be for the same task of recognizing animals from camera trap images, with each distribution corresponding to a different set of locations [9, 47]. We posit that using one small OOD validation set $D_{\text{val}}$ can guide the model to generalize better to various new OOD distributions. This validation set reflects the type of distribution shifts that the model might encounter. Performance on this set can serve as an effective proxy for performance on novel OOD distributions. Our experiments in Sec. 6 confirm that AutoFT can successfully leverage one small OOD validation set to improve generalization to novel distributions, outperforming state-of-the-art methods for robust fine-tuning.
@@
  Input Hyperparameter optimizer $\mathrm{HPO}$
  Input ID training data $D_{\text{tr}}$, OOD validation data $D_{\text{val}}$
  for $\phi\leftarrow\mathrm{HPO}.\mathrm{Sample}()$ do
@@
     $f_{ft}\leftarrow\text{LearnAlg}(D_{\text{tr}},\phi)$ // Fine-tune model
     $p\leftarrow\mathrm{Perf}(f_{ft},D_{\text{val}})$ // Evaluate
     $\mathrm{HPO}.\mathrm{Update}(\phi,p)$ // HPO Update
  end for
  $\phi^{*}\leftarrow\mathrm{HPO}.\mathrm{Best}()$ // Get best hyperparameters
  $\theta^{*}\leftarrow\text{LearnAlg}(D_{\text{tr}},\phi^{*})$ // Fine-tune final model
4.2 Hyperparameter Optimization Loop

### 07 Mar 2024 v2

# AutoFT: Learning an Objective for Robust Fine-Tuning

@@ Abstract
Foundation models encode rich representations that can be adapted to downstream tasks by fine-tuning. However, fine-tuning a model on one data distribution often degrades performance under distribution shifts. Current approaches to robust fine-tuning use hand-crafted regularization techniques to constrain the fine-tuning process towards the pretrained model. Yet, it is hard to specify how to adapt relevant characteristics of the foundation model during fine-tuning, as this depends on how the pre-training, fine-tuning, and test data distributions relate to each other. We propose AutoFT, a data-driven approach for robust fine-tuning. Given a task, AutoFT searches for a fine-tuning procedure that enhances out-of-distribution (OOD) generalization. Specifically, AutoFT uses bi-level optimization to search for an objective function and hyperparameters that maximize post-adaptation performance on a small OOD validation set. We evaluate AutoFT on nine natural distribution shifts. Our experiments show that AutoFT significantly improves generalization to OOD inputs, outperforming existing robust fine-tuning methods. Notably, AutoFT achieves a new state-of-the-art on the WILDS iWildCam and FMoW benchmarks, outperforming the previous best methods by $6.0\%$ and $1.5\%$, respectively.
Machine Learning, ICML
@@ 1 Introduction
Foundation models have emerged as a powerful tool in machine learning, demonstrating unprecedented performance across a wide variety of data distributions  (Radford et al., 2021b; Ilharco et al., 2021; Jia et al., 2021). By pre-training on large and diverse datasets, these models encode rich, common-sense knowledge that can be leveraged in downstream tasks to improve performance. However, while fine-tuning on additional task-specific data improves performance on the target distribution, it degrades performance on different distributions.
@@
Figure 1: Overview of AutoFT: a method for robustly fine-tuning foundation models. While fine-tuning with in-distribution (ID) data (blue), AutoFT searches for a fine-tuning objective that maximizes performance on a small out-of-distribution validation set (red). This validation set serves as a proxy for performance on different distributions (green and purple), allowing AutoFT to learn a robust fine-tuning procedure.
@@
Figure 2: A summary of our data assumptions and evaluation protocol. The standard approach is to optimize hyperparameters on a validation dataset drawn from the same distribution as the training data. In contrast, AutoFT employs a small out-of-distribution (OOD) validation set for hyperparameter optimization, enhancing the generalizability of the final model. We evaluate all fine-tuned models on data from unseen distribution shifts (green and purple).
@@
This issue has driven recent research on robust fine-tuning, which aims to produce an adapted model that achieves good performance under distribution shifts. Prior works have proposed various regularization techniques to preserve the prior knowledge embedded in the foundation model, such as ensembling models before and after adaptation (Wortsman et al., 2022b) or initially fitting only the last layer (Kumar et al., 2022a). However, these approaches propose hard-coded fine-tuning procedures, which may not fully capture the intricate relationship between foundation model priors and task-specific data during adaptation. Furthermore, this relationship can vary across fine-tuning tasks, and a one-size-fits-all approach may not account for these differences.
We introduce AutoFT, a novel approach for robust fine-tuning that aims to optimally balance between preserving foundation model priors and incorporating task-specific knowledge during adaptation. Our key insight is that we can learn the best adaptation strategy for a given fine-tuning task. AutoFT learns a fine-tuning objective and hyperparameters using a small out-of-distribution (OOD) validation set to produce models that are robust to distribution shifts. We then fine-tune the foundation model with the learned objective and hyperparameters and evaluate its robustness to new distribution shifts.
Our contributions are twofold. First, we learn the fine-tuning objective itself, parameterized by weight coefficients for several different loss functions and regularizers. This large search space gives AutoFT more granular control over adaptation. Second, AutoFT optimizes the fine-tuning objective and hyperparameters with respect to post-adaptation performance on a proxy distribution shift: an OOD validation set. Importantly, this validation set is drawn from a different distribution than the final evaluation datasets and is small, containing fewer than $1\%$ of the number of examples in the fine-tuning dataset. We illustrate the intuition behind our approach in Figure 1 and our data assumptions in Figure 2.
@@
We rigorously evaluate AutoFT on a wide array of real-world datasets and consider various types of distribution shifts, including subpopulation and domain shift. Our experiments show that a model fine-tuned with the objective learned by AutoFT generalizes better to previously unseen distribution shifts. With as few as $1000$ examples from a different distribution than that of the fine-tuning data, which is often readily available, AutoFT outperforms existing robust fine-tuning methods across all benchmarks. These gains in robustness are achieved with minimal additional compute, requiring at most $5\%$ more compute compared to standard fine-tuning. Among other results, AutoFT achieves new state-of-the-art performance on the challenging iWildCam and FMoW benchmarks (Beery et al., 2021; Koh et al., 2021; Christie et al., 2018), outperforming the prior best methods by $6.0\%$ and $1.5\%$, respectively.
@@ 2 Related Work
Transfer learning. While early research demonstrated that using features learned from pre-training on large datasets are effective for new tasks, transfer learning has evolved to optimize performance in settings with limited data  (Oquab et al., 2014; Yosinski et al., 2014; Sharif Razavian et al., 2014). Common transfer learning techniques include fine-tuning with regularization (Zhang et al., 2020; Xuhong et al., 2018; Lee et al., 2019a; Jiang et al., 2019; Li et al., 2020; Aghajanyan et al., 2020; Gouk et al., 2021; Shen et al., 2021; Karani et al., 2021) and selectively freezing pre-trained parameters (Kirkpatrick et al., 2017; Lee et al., 2019b; Guo et al., 2019; Ramasesh et al., 2020; Liu et al., 2021b; Royer & Lampert, 2020; Eastwood et al., 2021; Evci et al., 2022; Eastwood et al., 2022; Cohen et al., 2022; Touvron et al., 2022; Lee et al., 2022a; Kumar et al., 2022b). However, as pre-trained models are fine-tuned for a specific distribution, their effective robustness decreases at convergence (Andreassen et al., 2021). We introduce a transfer learning framework that preserves the robustness of the pre-trained model while adapting to new tasks.
AutoML and bi-level optimization. Our work leverages high-level ideas from the broader literature on meta-learning and hyperparameter optimization. Such methods have proposed to optimize different parts of the training pipeline, including general hyperparameters (Hutter et al., 2011; Bergstra & Bengio, 2012; Feurer et al., 2015; Li et al., 2017, 2018b), network architectures (Zoph & Le, 2016; Zoph et al., 2018; Liu et al., 2018; Real et al., 2019; Xu et al., 2019), augmentation policies (Cubuk et al., 2019; Lim et al., 2019; Hataya et al., 2020; Cubuk et al., 2020), optimizers (Bengio et al., 2013; Andrychowicz et al., 2016; Wichrowska et al., 2017; Metz et al., 2022; Chen et al., 2023), and objective functions (Yu et al., 2018; Kirsch et al., 2019; Oh et al., 2020; Bechtle et al., 2021). However, most of these works optimize for generalization within the training distribution and do not consider robustness to distribution shifts. Existing works that optimize a training procedure for OOD generalization consider a structured few-shot adaptation setting (Li et al., 2018a; Zhang et al., 2021), limiting their scalability to large datasets. Tian et al. (2023) learn layer-specific regularization constraints to improve OOD generalization; however, they do not learn a fine-tuning objective.
@@
Out-of-distribution generalization. Maintaining good performance on data that deviates from the training distribution is crucial in many real-world applications, where models may face data from unfamiliar environments (Hendrycks & Dietterich, 2019; Geirhos et al., 2020; Gulrajani & Lopez-Paz, 2020; Koh et al., 2021). Numerous studies have investigated how to ensure robustness to various distribution shifts (Tzeng et al., 2014; Byrd & Lipton, 2019; Hendrycks et al., 2019; Arjovsky et al., 2019; Salman et al., 2020; Liu et al., 2021a; Wiles et al., 2021; Andreassen et al., 2021; Creager et al., 2021; Lee et al., 2022b). Some works have shown that despite the myriad of ways in which data distributions can change, naturally occurring distribution shifts have a surprisingly predictable effect on model performance (Taori et al., 2020; Miller et al., 2021; Baek et al., 2022), suggesting that it may be possible to learn how to be robust to unseen natural distribution shifts. Similar to our work, Goyal et al. (2022b) meta-learn an objective, but for test-time adaptation instead of fine-tuning. Benchmarks for OOD generalization, such as DomainBed (Gulrajani & Lopez-Paz, 2020) and WILDS (Koh et al., 2021), have discussed data assumptions for dealing with distribution shifts, and suggested that an OOD validation set can be a useful signal for hyperparameter optimization and model selection. AutoFT leverages this signal much more directly than existing works by learning an objective for fine-tuning.
@@
Robust fine-tuning. Foundation models trained on massive datasets encode a broad range of general knowledge, enabling robust performance across various data distributions, including OOD scenarios (Radford et al., 2021b; Bommasani et al., 2021). While in principle, foundation models should serve as a useful prior for further fine-tuning, empirical evidence shows that fine-tuning foundation models on a new task often leads to a significant drop in OOD performance (Andreassen et al., 2021). Recent works have proposed modifications to the basic fine-tuning procedure to improve OOD generalization (Wortsman et al., 2022b; Kumar et al., 2022a; Wortsman et al., 2022a; Goyal et al., 2022a; Mukhoti et al., 2023). Instead of hand-designing a regularization technique, we propose a data-driven approach to learn a more nuanced fine-tuning procedure. In fact, some prior works (Xuhong et al., 2018; Lee et al., 2022a; Goyal et al., 2022a) can be seen as special cases of AutoFT since our hyperparameter search space encompasses these fine-tuning algorithms. Our experiments in Section 6 demonstrate that AutoFT consistently outperforms prior robust fine-tuning methods on OOD data.
3 Background: Hyperparameter Optimization
We first formalize hyperparameter optimization as a means for searching the space of fine-tuning procedures in Section 4. Hyperparameters are predefined properties of the learning algorithm which are not learned during training, such as network architecture, learning rate, and regularization strength. A good choice of hyperparameters can significantly improve model performance, yet optimal hyperparameters vary, depending on the data distribution and evaluation metrics.
@@ 3 Background: Hyperparameter Optimization
Formally, we denote the learning algorithm as LearnAlg and its hyperparameters as $\phi\in\Phi$, where $\Phi$ is the hyperparameter space. We denote the training and validation datasets as $D_{\text{tr}}$ and $D_{\text{val}}$, respectively. These datasets are disjoint and typically drawn from the same distribution. The learning algorithm LearnAlg produces a model by training on $D_{\text{tr}}$ with hyperparameters $\phi$. We denote the resulting model as $\text{LearnAlg}(\phi,D_{\text{tr}})$. Hyperparameter optimization finds hyperparameters that maximize some performance metric $\mathrm{Perf}(f,D_{\text{val}})$ which depends on the model $f$ and the validation dataset $D_{\text{val}}$. Examples of performance metrics include top-1 accuracy, macro F1 score, and worst-region accuracy. We formalize hyperparameter optimization as
@@
Hyperparameter optimization methods typically start with randomly initialized model parameters and use a validation set $D_{\text{val}}$, drawn from the same distribution as the training data, to adjust hyperparameters. The problem of robust fine-tuning, however, begins with pre-trained model parameters and aims to achieve high performance on test data that diverges from the fine-tuning distribution. In the next section, we describe how we leverage hyperparameter optimization to learn a robust adaptation procedure.
4 AutoFT: Robust Fine-Tuning with a Learned Objective
In this section, we present AutoFT, a data-driven approach for robust fine-tuning. Given a foundation model and task-specific data, AutoFT uses bi-level optimization to learn a fine-tuning objective and optimal hyperparameters. There are two key components in AutoFT. First, learning the fine-tuning objective enables a more precise adaptation to task-specific data. Second, AutoFT optimizes the objective and hyperparameters for performance under a proxy distribution shift, which enhances OOD generalization.
@@ 4.1 Problem Setting: Robust Fine-Tuning
Data assumptions. We assume access to two datasets: (1) a large fine-tuning dataset $D_{\text{tr}}$ from the training distribution $\mathcal{P}_{\text{tr}}$, and (2) a small held-out OOD validation set $D_{\text{val}}$ from a different distribution $\mathcal{P}_{\text{val}}$. The validation set $D_{\text{val}}$ is much smaller than $D_{\text{tr}}$, containing as few as $1000$ examples, and is only used for outer-level optimization, not for fine-tuning. We then fine-tune on $D_{\text{tr}}$ with the found hyperparameters and test the fine-tuned model on several OOD distributions $\mathcal{P}_{\text{ood}}^{1},\mathcal{P}_{\text{ood}}^{2},\dots$. Note that these test distributions are distinct from both $\mathcal{P}_{\text{tr}}$ and $\mathcal{P}_{\text{val}}$, and are never shown to the model before final evaluation. Following prior works on robust fine-tuning, all distributions are assumed to be related to the same task. We consider domain and subpopulation shifts arising from differences in data collection. For instance, in animal recognition from camera trap images, each distribution may represent different locations (Beery et al., 2021; Koh et al., 2021). Such data is often readily available, as shown in the WILDS and DomainBed benchmarks (Koh et al., 2021; Gulrajani & Lopez-Paz, 2020), and can be constructed by partitioning by time or geographical location.
Notation and concrete setup. Let $f$ denote a pre-trained foundation model with parameters $\theta$, which we adapt by fine-tuning on $D_{\text{tr}}$. We denote the fine-tuning algorithm as $\text{LearnAlg}(\phi,D_{\text{tr}})$, where $\phi$ specifies both the fine-tuning objective and other hyperparameters such as learning rate. In our experiments, the performance metric $\mathrm{Perf}$ follows the standards in each evaluation setting and is either top-1 accuracy, worst-region accuracy, or macro F1.
@@
0:  Hyperparameter Optimizer (HPO)
0:  ID Training Data $D_{\text{tr}}$, OOD Validation Data $D_{\text{val}}$
[…]
2:     $f_{ft}\leftarrow\text{LearnAlg}(D_{\text{tr}},\mathcal{L}_{\phi})$ // Short fine-tuning run with sampled loss
3:     $p\leftarrow\mathrm{Perf}(f_{ft},D_{\text{val}})$ // Evaluation on OOD Val set
4:  end for
5:  $\phi^{*}\leftarrow\mathrm{HPO}.\mathrm{Best}()$ // Get best hyperparameters
@@
6:  $f^{*}\leftarrow\text{LearnAlg}(D_{\text{tr}},\mathcal{L}_{\phi^{*}})$ // Fine-tune with loss $\mathcal{L}_{\phi^{*}}$
7:  Return $f^{*}$
4.2 Learning a Fine-Tuning Objective via Bi-Level Optimization